# Learning to Predict Structural Vibrations

### v3

@@ 3.2 Decoder $\Psi$
To incorporate physics-based contraints and take advantage of the larger amount of available data, we employ a UNet to predict the velocity fields, $\mathbf{V}(f)$. From this, we derive the frequency response, $\mathcal{F}(f)$, using Equation 1. A frequency query, introduced via a film layer after the encoder, enables frequency-specific predictions. The downside of frequency specific predictions and evaluating the decoder for all 300 frequency queries is increased computational cost and memory demands, impacting training and inference. To interpolate between a query-based decoder and predicting all frequency response values at once, we predict a neighborhood around each query. Reducing the necessary computations by a factor of five, predictions for five velocity fields per query are generated. The training loss is set as $L_{total}=\alpha L_{\mathbf{V}}+(1-\alpha)L_{F}$, where $L_{\mathbf{V}}$ and $L_{F}$ represent the MSE on the log-transformed, normalized velocity field and normalized frequency response respectively. $\alpha$ is set to 0.9 focusing on $L_{\mathbf{V}}$ (Ablation results for $\alpha$ in Appendix D).
@@ 3.3 Baseline Methods
We further report baseline results on the following alternative methods: A $k$-Nearest Neighbors regressor, that finds the nearest neighbors in the latent space of an autoencoder. DeepONet [11], with a RN18 as backbone and a MLP to encode the query frequencies as a branch net. Two architectures based on Fourier Neural Operators [15]. One employing a FNO as a replacement for the query-based decoder based on RN18 features. The second directly takes the input geometry and is trained to map it to the velocity fields.
@@ 4 Experiments
$k$-NN - - 0.63 21.50 0.45 8.7 0.83 31.31 0.68 18.0
RN18 + FNO - - 0.42 10.76 0.34 5.6 0.33 12.76 0.37 6.2
DeepONet $\checkmark$ - 0.49 16.91 0.48 5.4 0.47 20.56 0.64 6.8
FNO (velocity field) - $\checkmark$ 0.47 13.10 0.36 6.3 0.34 13.31 0.38 6.2
Grid-RN18 - - 0.44 13.29 0.36 5.4 0.28 12.11 0.32 5.3
FQO-RN18 (Q1) $\checkmark$ - 0.32 10.70 0.17 5.3 0.22 10.37 0.16 4.8
FQO-ViT (Q2) $\checkmark$ - 0.68 20.96 0.54 7.1 0.53 22.23 0.54 9.3
Grid-UNet - $\checkmark$ 0.19 7.57 0.24 2.7 0.15 6.93 0.17 3.5
FQO-UNet $\checkmark$ $\checkmark$ 0.09 3.90 0.08 1.8 0.09 4.94 0.07 2.5
We train the architecture variations and baseline methods on the vibrating-plates dataset (see Table 1). To assess which architecture aspects described in Section 3 are beneficial, we perform the following comparisons. Regarding (Q1 - frequency-query approach), the Frequency-Query Operator variations consistently yield better predictions than equivalent grid-based methods, where responses for all frequencies are predicted at once: The $\mathcal{E}_{\text{MSE}}$ and the $\mathcal{E}_{\text{EMD}}$ are lower, more peaks are reproduced, and the peak positions are more precise. Regarding (Q3 - velocity field prediction), predicting the velocity fields and then transforming them to the frequency response leads to better results than directly predicting the frequency response. Specifically, the UNet based architectures strongly outperform all alternatives, which we attribute to the richer training data of velocity fields. Regarding (Q2 - ViT encoder), the ViT encoder leads to worse results than the cnn-based encoders.
All evaluated baseline methods achieve comparatively worse results than our Frequency-Query Operator methods. Despite using the same RN18 geometry encoder as FQO-RN18, DeepONet [11] performs worse. We assume that this is due to its approach of incorporating frequency information through a single weighted summation, which limits the model’s expressivity [37]. In contrast, FQO-RN18 introduces the queried frequency earlier into the model. Two Fourier Neural Operator [15, FNO] baseline methods are evaluated: the first, RN18 + FNO, which substitutes the query-based decoder with an FNO decoder, underperforms compared to FQO-RN18 on both datasets. The second FNO baseline, trained directly to predict velocity fields, yields poorer results despite having access to richer training data.
Results for the G-5000 setting are similar to the V-5000 setting with some difference in individual metrics but no clear trend. The lack of clear differences is surprising because the space of plates in the G-5000 setting is a superset of the V-5000 space. One reason might be the average number of peaks in the frequency response: the plates in G-5000 are on average smaller and because of this stiffer, leading to fewer peaks (on average 3.9 in G-5000 and 5.9 in V-5000). This interpretation is supported by the fact that the average error becomes higher with increasing frequency (Figure 5f) and thus increasing peak density (Figure 3d).
@@
Figure 5: Results. (b) to (d) show the velocity field at one frequency and prediction for the plate geometry in (a) from FQO-UNet. (e) shows the test MSE for training two methods with reduced numbers of samples from V-5000. (f) shows the MSE for the FQO-UNet for different frequencies.
@@
To quantify to which degree features learned on a subset of the design space transfer to a different subset, the V-5000 setting is split into two equally-sized parts based on the number of mesh elements that are part of a beading. The ”more beadings” set contains only 5.1 peaks on average because the plates are stiffened by the beadings, compared to 6.7 peaks on average for the ”less beadings” set. The training on plates with less beadings leads to a smaller drop in prediction quality (see Table 2). This indicates that training on data with more complex frequency responses might be more efficient.
Table 2: Transfer learning performance: We split V-5000 into two halves based on amount of beadings and evaluate transfer learning performance across these splits: training subset $\mapsto$ test subset. The gray rows denote validation results on the original subset that has been used for training.
@@
FQO-UNet 0.41 9.64 0.14 6.4 0.55 9.44 0.21 6.1
(origin) 0.16 6.03 0.13 3.2 0.18 6.08 0.10 3.2
@@ 5 Related Work
While research on surrogate models for the spatio-temporal evolution of vector fields is fairly common, directly predicting frequency responses through neural networks is an understudied problem. A general CNN architecture is applied in [38] to calibrate the parameters of an analytical model for a composite column on a shake table. The data includes spectrograms representing the structural response in time-frequency domain. The frequency-domain response of acoustic metamaterials is considered in a material design task by conditional generative adversarial networks or reinforcement learning [39, 40, 41]. The frequency response of a multi-mass oscillator is predicted with transformer-based methods [42]. Within the context of aeroacoustics, the propagation of a two-dimensional acoustic wave while considering sound-scattering obstacles is predicted in time-domain by a CNN [43, 44]. A review of machine learning in acoustics is given by [45]. Concerning benchmarks in acoustics, with [46], several acoustic benchmarks for numerical methods are available. However, these benchmarks do not systematically vary input geometries, making them not directly applicable to data-driven models.
@@
A promising avenue of research for incorporating inductive biases for physical models has been operator learning [11, 15, 64, 37, 65]. Operator learning structures neural networks such that instead of directly mapping from discrete input to output space, the neural network produces a function that can be evaluated at real values instead of a discrete grid. DeepONet [11] implements operator learning by taking the value at which it is evaluated as an input and processes this value in a separate branch. Fourier Neural Operators [15] use a point-wise mapping to a latent space which is processed through a sequence of individual layers in Fourier space before being projected to the output space.
@@ 6 Conclusion
We introduced the problem of predicting structural vibrations and associated frequency response functions of mechanical systems. Unlike other benchmarks for PDE deep learning surrogate models, this task necessitates predicting a steady-state solution that remains constant over time, but varies across different excitation frequencies. To this end, we created the vibrating-plates dataset and benchmark and provide reference scores for several methods. Our Frequency-Query Operator method addresses the benchmark and achieves better results than the DeepONet and FNO baselines. Surrogate models as shown in this work can greatly accelerate the prediction of physical quantities over the finite element method: Our models achieved a speed-up of around 4 to 6 orders of magnitude (see Appendix C). This efficiency, however, depends on the availability of pre-generated training data and necessitates model training. Our results underscore the necessity of selecting an appropriate neural network architecture for accurate frequency response predictions. Specifically, query-based approaches and the indirect prediction of a mean frequency response through predicted field quantities led to better results. Broadly, our experiments demonstrate that steady-state vibrations can be learned with data-driven methods.
Limitations.
Our dataset and method serve as an initial step in the development of surrogate models for vibration prediction. The dataset focuses on plates, a common geometric primitive used in a great number of applications. However, many structures beyond plates exist, involving curved shells, multi-component geometries and complex material parameters. While our approach remains applicable in such cases, appropriate encoders and decoders would need to be designed. As more complex geometries incur high computational costs of FEM simulations, a key question is how to enhance sample-efficiency further. A further limitation is the manufacturability of the considered beading patterns. The plate beadings could in principle be manufactured by stamping plates, but would require specifically designed stamps.
@@ Acknowledgements.
This research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), Project number 501927736, within the DFG Priority Programme ’SPP 2353: Daring More Intelligence - Design Assistants in Mechanics and Dynamics’. This support is highly appreciated. Additionally, we thank Harikrishnan K. Sreekumar for technical support with the FEM software.
@@ A.1 The Mechanical Model
This equation is combined with a disc formulation for in-plane loads in order to receive a shell formulation for the mechanical description of arbitrarily formed moderately thin structures considering in-plane and transverse loads. The plate part is the dominating and important part for resolving bending waves. In the equation, $u_{z}$ denotes the normal displacement of the plate structure as degree of freedom of interest. $B$ represents the bending stiffness, $\rho_{s}$ the density, $h$ the thickness, $G$ the shear modulus and $I$ the moment of inertia. The angular frequency $\omega$ is defined as $\omega=2\pi f$. The right hand-side excitation $p_{l}$ describes an applied pressure load, which is converted to point forces through integration. As boundary conditions we apply homogenious dirichlet boundary conditions, i.e. $u_{z}(x)=0$ on the boundary. The equation is transformed into a weak integral formulation by weighted residuals, discretized using finite elements and integrated numerically. In particular, we use triangular shell elements with 3 nodes and linear ansatz functions. The integration delivers the sparse linear system of equations. This linear system is solved using the direct solver MUMPS [71] with a specialized FEM implementation for acoustics [24]. The discretization is chosen, such that the bending waves are resolved with a minimum of 10 nodes. The bending wave length $\lambda_{B}$ of a plate can be calculated by
@@
where $E$ is the Young’s modulus, $t$ the thickness, $\nu$ the Poisson ratio and $\rho$ the density of the plate. The final discretization is set to $121\times 81$, which is sufficient for convergence.
@@ A.2 Datasets
The exact physical setting and variation of our mechanical model to form the V-5000 and G-5000 datasets are given in Table 4 and Table 3. Both dataset settings contain 6000 samples, each consisting of a plate geometry with associated physical and material parameters and the computed velocity fields $\mathbf{V}(f)$ and frequency response $\mathcal{F}(f)$ for frequencies $f$ 1 to 300 $Hz$. 1000 samples from the 6000 samples are selected as a test dataset and not considered during neural network training or validation. Computing a single sample out of the 6000 samples takes 2 minutes and 19 seconds on a machine with a 2 Ghz CPU (20 physical cores).
Table 3: Dataset settings. Width is the width of lines and ellipses in mm. Geometry (geom.) involves plate size, thickness and material.
@@
Setting Geom. Lines Ellipses Width Train Test
@@
Table 4: Geometry and physical parameters for V-5000 and G-5000 datasets.
@@
For the G-5000 dataset, the geometry and material parameters are varied. The effect of two of those material parameters is visualized in Figure 6. Increasing the damping reduces amplitudes at resonance peaks but does not shift the overall form of the frequency response. Increasing the thickness increases the overall stiffness shifts resonance peaks towards higher frequencies in a less regular manner.

### v4

@@ 3.2 Decoder $\Psi$
To incorporate physics-based contraints and take advantage of the larger amount of available data, we employ a UNet to predict the velocity fields, $\mathbf{V}(f)$. From $\mathbf{V}(f)$, we derive the frequency response $\mathcal{F}(f)$ (analogous to Equation 1). A frequency query, introduced via a FiLM layer after the encoder, enables frequency-specific predictions. To reduce the memory and computation demands per geometry during training, we select a random subset of $k$ frequency queries per geometry in a batch, with $k<300$. If not otherwise specified, $k$ is set to 50.
@@ 3.3 Baseline Methods
We further report baseline results on the following alternative methods: A $k$-Nearest Neighbors regressor, that finds the nearest neighbors in the latent space of an autoencoder. DeepONet [11], with a RN18 as backbone and a MLP to encode the query frequencies as a branch net. Two architectures based on Fourier Neural Operators [15]. One employing an FNO as a replacement for the query-based decoder based on RN18 features. The second directly takes the input geometry and is trained to map it to the velocity fields.
3.4 Training
All methods are trained in a data-driven fashion for 500 epochs on the training dataset of 5000 samples. 500 samples from the training dataset are excluded and employed for validation. We report evaluation results on the previously unseen test set consisting of 1000 additional samples.
For methods that predict $\mathbf{V}(f)$, i.e. UNet based methods and the FNO variation, the training loss is set to $L_{\mathbf{V}}$ where $L_{\mathbf{V}}$ represents the MSE on the log-transformed, normalized squared velocity field (Ablation on loss function in Appendix D). For methods that directly predict $\mathcal{F}$, the loss is set to $L_{F}$, the MSE on the normalized frequency response. Choosing the log-transformed quantities enables the loss to be sensitive to errors outside of resonance frequencies. Otherwise, such errors would have little influence on the total loss, as their magnitude is much lower. See Appendix C for further details on the architectures and training procedure.
@@ 4 Experiments
$k$-NN - - 0.63 21.50 0.45 8.7 0.88 32.48 0.68 21.0
RN18 + FNO - - 0.42 10.76 0.34 5.6 0.28 14.12 0.21 6.1
DeepONet $\checkmark$ - 0.49 16.91 0.48 5.4 0.44 23.05 0.57 9.9
FNO (velocity field) - $\checkmark$ 0.47 13.10 0.36 6.3 0.49 21.16 0.39 10.7
Grid-RN18 - - 0.44 13.29 0.36 5.4 0.30 14.95 0.26 6.5
FQO-RN18 (Q1) $\checkmark$ - 0.32 10.70 0.17 5.3 0.24 13.51 0.13 5.1
FQO-ViT (Q2) $\checkmark$ - 0.68 20.96 0.54 7.1 0.52 24.34 0.49 11.5
Grid-UNet - $\checkmark$ 0.19 7.57 0.24 2.7 0.17 9.41 0.14 4.6
FQO-UNet $\checkmark$ $\checkmark$ 0.08 4.24 0.07 1.7 0.11 7.47 0.08 3.1
We train the architecture variations and baseline methods on the Vibrating Plates dataset (see Table 1). To assess which architecture aspects described in Section 3 are beneficial, we perform the following comparisons. Regarding Q1 (frequency-query approach), the Frequency-Query Operator variations consistently yield better predictions than equivalent grid-based methods, where responses for all frequencies are predicted at once: The $\mathcal{E}_{\text{MSE}}$ and the $\mathcal{E}_{\text{EMD}}$ are lower, more peaks are reproduced, and the peak positions are more precise. Regarding Q3 (velocity field prediction), predicting the velocity fields and then transforming them to the frequency response leads to better results than directly predicting the frequency response. Specifically, the UNet based architectures strongly outperform all alternatives, which we attribute to the richer training data of velocity fields. Regarding Q2, the ViT encoder leads to worse results than the CNN-based encoders.
All evaluated baseline methods achieve comparatively worse results than our proposed methods. Despite using the same RN18 geometry encoder as FQO-RN18, DeepONet [11] performs worse. We assume that this is due to incorporating frequency information through a single weighted summation, which limits the model’s expressivity [37]. In contrast, FQO-RN18 introduces the queried frequency earlier into the model. Two Fourier Neural Operator [15] baseline methods are evaluated: the first, RN18 + FNO, which substitutes the query-based decoder with an FNO decoder, underperforms compared to FQO-RN18 on both datasets. The second FNO baseline, trained directly to predict velocity fields, yields poorer results.
Results for the G-5000 setting are slightly worse than for the V-5000 setting. The difference is surprising small considering the seven additional varied parameters in the G-5000 setting. One reason might be the average number of peaks in the frequency response: the plates in G-5000 are on average smaller and because of this stiffer, leading to fewer peaks (on average 3.9 in G-5000 and 5.9 in V-5000). This interpretation is supported by the fact that the average error becomes higher with increasing frequency and thus increasing peak density (Figure 3d).
@@
Figure 5: Results. (b) to (d) show the velocity field at one frequency and prediction for the plate geometry in (a) from FQO-UNet. (e) shows the test MSE for training two methods with reduced numbers of samples from V-5000. (f) shows effects of different data generation strategies. The blue line is an isoconture for a fixed compute budget of 150,000 data points, with varying number of frequencies per plate geometry. The green star represents using a larger dataset at 15 frequencies per plate (half of V-5000). The red cross represents a model trained on V-5000. Training with fewer frequencies per plate is more efficient.
@@
To quantify to which degree features learned on a subset of the design space transfer to a different subset, the V-5000 setting is split into two equally-sized parts based on the number of mesh elements that are part of a beading. The ”more beadings” set contains only 5.1 peaks on average because the plates are stiffened by the beadings, compared to 6.7 peaks on average for the ”less beadings” set. The training on plates with less beadings leads to a smaller drop in prediction quality (see Table 2). This indicates that training on data with more complex frequency responses might be more efficient. In addition, we train a single model on both G-5000 and V-5000. Performance increases, indicating that training can benefit from training with data based on similar mechanical models (Table 3).
Table 2: Transfer learning performance: We split V-5000 into two halves based on amount of beadings and evaluate transfer learning performance across these splits: training subset $\mapsto$ test subset. The gray rows denote test results on the original subset that has been used for training.
@@
FQO-UNet 0.39 11.17 0.21 5.6 0.54 12.02 0.25 5.5
(origin) 0.18 8.68 0.19 2.6 0.17 7.83 0.13 3.0
Table 3: A FQO-UNet is trained in parallel on batches from V-5000 and G-5000 and evaluated on the G-5000 test set. Performance increases in all metrics.
$\mathcal{E}_{\text{MSE}}$ $\mathcal{E}_{\text{EMD}}$ $\mathcal{E}_{\text{PEAKS}}$ $\mathcal{E}_{\text{F}}$
G-5000 0.111 7.47 0.079 3.1
G-5000 + V-5000 0.093 6.97 0.071 2.9
@@
We further investigate the optimal ratio of numbers of frequencies per geometry and total number of geometries, by generating an additional dataset in the V-5000 setting consisting of 50,000 plate geometries but with only 15 frequency evaluations per geometry. These frequencies are uniformly spaced with a random starting frequency. Reducing the frequencies per geometry drastically increases the data efficiency of our method. With a tenth of data points compared to our original dataset, the MSE metric approaches the original value (Figure 5f, quantitative results in Appendix E.1).
Design optimization.
We investigate the potential of our FQO-UNet to be used for optimizing a beading pattern for reduced vibrations in a specified frequency range. Following the approach described in [38], to generate plates with reduced vibrations, a diffusion model trained to generate novel beading patterns is combined with gradient information from our FQO-UNet as follows: A gradient on the pixels of the input beading pattern is obtained by passing a beading pattern through the network, computing the sum of the predicted frequency response as a loss and then performing backpropagation to the input beading pattern. This gradient is then used to guide the diffusion model to generate beading patterns with reduced vibrations. We optimize beading patterns to reduce vibrations between 100 and 200 Hz using the FQO-UNet trained on the V-5000 dataset (Figure 6). Resulting plates have a lower mean frequency response in the targeted range than any plate in the training dataset.
Figure 6: Design optimization. Exemplary generation result with lowest mean response between $100$ Hz and $200$ Hz out of 32 generations (left, mean response below). Plate with lowest response out of all 5000 training examples from V-5000 (middle left). Comparison of responses from left plates (middle right). Responses from 16 generated plates (right).
@@ 5 Related Work
While research on surrogate models for the spatio-temporal evolution of vector fields is fairly common [39, 40, 41], directly predicting frequency responses through neural networks is an understudied problem. A general CNN architecture is applied in [42] to calibrate the parameters of an analytical model for a composite column on a shake table. The data includes spectrograms representing the structural response in time-frequency domain. The frequency-domain response of acoustic metamaterials is considered in a material design task by conditional generative adversarial networks or reinforcement learning [43, 44, 45]. The frequency response of a multi-mass oscillator is predicted with transformer-based methods [46]. Within the context of aeroacoustics, the propagation of a two-dimensional acoustic wave while considering sound-scattering obstacles is predicted in time-domain by a CNN [47, 48]. A review of machine learning in acoustics is given by [49]. Several acoustic benchmarks for numerical methods are available [50], however, these benchmarks do not systematically vary input geometries, making them not directly applicable to data-driven models.
@@
A promising avenue of research for incorporating inductive biases for physical models has been operator learning [11, 15, 66, 37, 41]. Operator learning structures neural networks such that they implement a function that can be evaluated at real values instead of a fixed discrete grid. DeepONet [11] implements operator learning by taking the value at which it is evaluated as an input and processes this value in a separate branch. Fourier Neural Operators [15] use a point-wise mapping to a latent space which is processed through a sequence of individual layers in Fourier space before being projected to the output space.
@@ 6 Conclusion
We introduced the problem of predicting structural vibrations and associated frequency response functions of mechanical systems. Unlike other benchmarks for deep learning surrogate models, this task necessitates predicting a steady-state solution that remains constant over time, but varies across different excitation frequencies. To this end, we created the Vibrating Plates dataset and benchmark and provide reference scores for several methods. Our Frequency-Query Operator method addresses the benchmark and achieves better results than the DeepONet and FNO baselines. We find that query-based approaches and the indirect prediction of a mean frequency response through predicted field quantities lead to better results. Surrogate models as shown in this work can greatly accelerate the prediction of physical quantities over the finite element method: Our models achieved a speed-up of around 4 to 6 orders of magnitude (see Appendix C), which makes tasks such as design optimization feasible. This efficiency, however, depends on the availability of enough pre-generated training data and requires model training. We further investigated effects of changing the composition of the training dataset and found that using less frequencies per plate and more different plates positively impacts prediction accuracy.
Limitations and future work.
Our dataset and method serve as an initial step in the development of surrogate models for vibration prediction. The dataset focuses on plates, a common geometric primitive used in a great number of applications. However, many structures beyond plates exist, involving curved shells, multi-component geometries and complex material parameters. While some results from our study might transfer to these cases, more flexible architectures, able to deal with 3D data, would be needed. Different mechanical models, might also produce more complex frequency responses with e.g. more closely spaced modes, making the prediction task more challenging. As more complex geometries incur higher computational costs of FEM simulations, key questions are how to enhance sample-efficiency further, for example through transfer learning. A further limitation is the manufacturability of the considered beading patterns. The plate beadings could in principle be manufactured by deep drawing of sheet metal, but would require specifically designed stamps.
@@ Acknowledgements.
This research is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), project number 501927736, within the DFG Priority Programme 2353: Daring More Intelligence - Design Assistants in Mechanics and Dynamics’. The authors gratefully acknowledge the computing time made available to them on the high-performance computers HLRN-IV at GWDG at the NHR Centers NHR@Göttingen. These centers are jointly supported by the German Federal Ministry of Education and Research and the German state governments participating in the NHR (www.nhr-verein.de/unsere-partner).
@@ A.1 The Mechanical Model
This equation is combined with a disc formulation for in-plane loads in order to receive a shell formulation for the mechanical description of arbitrarily formed moderately thin structures considering in-plane and transverse loads. The plate part is the dominating and important part for resolving bending waves. In the equation, $u_{z}$ denotes the normal displacement of the plate structure as degree of freedom of interest. $B$ represents the bending stiffness, $\rho_{s}$ the density, $h$ the thickness, $G$ the shear modulus and $I$ the moment of inertia. The angular frequency $\omega$ is defined as $\omega=2\pi f$. The right hand-side excitation $p_{l}$ describes an applied pressure load, which is converted to point forces through integration. As boundary conditions we apply homogeneous dirichlet boundary conditions, i.e. $u_{z}(x)=0$ on the boundary and include a rotational stiffness at the boundary to model different boundary conditions, ranging from free rotating to clamped plates. The equation is transformed into a weak integral formulation by weighted residuals, discretized using finite elements and integrated numerically. In particular, we use triangular shell elements with 3 nodes and linear ansatz functions. The integration delivers the sparse linear system of equations. This linear system is solved using the direct solver MUMPS [72] with a specialized FEM implementation for acoustics [24]. The discretization is chosen, such that the bending waves are resolved with a minimum of 10 nodes. The bending wave length $\lambda_{B}$ of a plate can be calculated by
@@
where $E$ is the Young’s modulus, $t$ the thickness, $\nu$ the Poisson ratio and $\rho$ the density of the plate. The final discretization is set to $181\times 121$ for G-5000 and $121\times 81$ for V-5000, which is sufficient for convergence.
@@ A.2 Datasets
[…]
Table 4: Dataset settings. Width is the width of lines and ellipses in mm. Properties. (prop.) involves plate size, thickness, material, boundary and loading properties.
@@
Setting Prop. Lines Ellipses Width Train Test
@@
Table 5: Geometry and material parameters for V-5000 and G-5000 datasets.
@@
Table 6: Loading and boundary condition parameters for V-5000 and G-5000 datasets.
Loading (Point force) Boundary condition (rot. stiffness)
x-position y-position $c_{ry}$/$c_{rx}$
V-5000 0.360.360.36\,0.36m 0.2250.2250.225\,0.225m 0.00.00.0\,0.0 Nm
G-5000 $0.18$ 0.720.720.72\,0.72 m $0.12$ 0.480.480.48\,0.48m $0.0$ 100100100\,100 Nm
@@
For the G-5000 dataset, the geometry, material, boundary condition and loading parameters are varied. The effect of two of the material parameters is visualized in Figure 7. Increasing the damping reduces amplitudes at resonance peaks but does not shift the overall form of the frequency response. Increasing the thickness increases the overall stiffness and shifts resonance peaks towards higher frequencies in a less regular manner. For boundary condition variation, we include one rotational stiffness parameter, which models the rotational stiffness along the x-axis at the lower and upper edge and along the y-axis at the left and right edge. The rotational stiffness is added at the respective rotational degree of freedom at the boundaries and varied as given in Table 6.